# Sharpness-Aware Minimization Can Hallucinate Minimizers

### v1

@@ 2.1 Simplified existence proof with isolated maximizers
We provide a brief sketch of the argument, with full details deferred to Appendix A.1. Figure 2 illustrates the key idea of the construction.
[…]
Next, consider $g(x)=\|x-x_{\star}\|^{2}$ and define
$x_{h}\in\operatorname*{arg\,max}_{x\in C_{\varepsilon}}g(x)=\operatorname*{arg\,max}_{x\in\partial C_{\varepsilon}}g(x).$
In words, if $x_{h}$ maximizes $g$ over $C_{\varepsilon}$, then it must lie on the boundary $\partial C_{\varepsilon}$. Since $\partial C_{\varepsilon}$ coincides with $\{x:f(x)=f(x^{\bullet})-\varepsilon\}$ near $x_{h}$, we apply the method of Lagrange multipliers to $x_{h}\in\operatorname*{arg\,max}_{x\in\partial C_{\varepsilon}}g(x)$ to obtain
$2(x_{\star}-x_{h})=\nabla g(x_{h})=\lambda\nabla f(x_{h}).$
The fact that $\lambda>0$ follows from the observation that $\nabla f(x_{h})$ points “toward” both $x_{\star}$ and $x^{\bullet}$, as illustrated in Figure 2. Finally, setting $\rho=\lambda\|\nabla f(x_{h})\|/2$ yields
$x_{h}+\rho\tfrac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{\star},$
and hence $f^{\mathrm{SAM}}(x_{h})=f(x_{\star})$. Thus, $x_{h}$ is a (global) minimizer of $f^{\mathrm{SAM}}$. ∎
@@
Importantly, the existence of a hallucinated minimizer also holds when $x_{\star}$ is a local minimizer, provided that $f$ has a locally Lipschitz gradient. See Appendix A.2 for details.
@@
The proof of Theorem 2.1 reveals the conditions under which a hallucinated minimizer is likely to arise. Figure 2 illustrates the core idea: given a minimizer $x_{\star}$, $x_{h}$ satisfies $x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}$. This means that at $x_{h}$, the gradient points directly toward $x_{\star}$. However, near a minimizer, the gradient typically points outward, making it difficult to identify such $x_{h}$ in its immediate neighborhood. To resolve this, Theorem 2.1 requires the presence of a nearby maximizer, which allows the gradient to align in the desired direction and thus necessitates nonconvexity of the objective.
The proof also shows that the perturbation radius $\rho$ must exceed the distance between minimizer and the maximizer, since $x_{h}$ is the farthest point from $x_{\star}$ on the superlevel set $C_{\varepsilon}$. Because $x_{h}\in C_{\varepsilon}$, a hallucinated minimizer is located near the maximizer. This implies that hallucinated minimizers are typically associated with high loss values.
Taken together, these observations suggest that hallucinated minimizers generally arise in nonconvex objectives containing both local maximizers and minimizers, and that they tend to occur near a local maximizer within a high-loss region.
2.2 Existence of hallucinated minimizers for neural networks
Theorem 2.1 assumes that the local maximizer is isolated. We now relax this assumption, since in neural networks, maximizers (like minimizers) often occur as sets.
To this end, we leverage the real-analyticity of neural networks, an assumption that holds when training on a finite dataset (empirical loss) with real-analytic activation functions. We note that all commonly used activation functions are real-analytic, except for ReLU. Technically, real-analyticity affords us the Łojasiewicz inequality, which we use to rule out certain pathological cases.
Notably, our result does not rely on restrictive but common structural assumptions on the loss function, such as global smoothness or a quadratic form.
Definition.
@@ 2.2 Existence of hallucinated minimizers for neural networks
Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function. A nonempty connected set $X^{\bullet}\subseteq\mathbb{R}^{d}$ is a local maximizer set of $f$ if there exists $\delta>0$ such that $X^{\bullet}$ is the maximizer set over its $\delta$-neighborhood. In other words,
$X^{\bullet}=\operatorname*{arg\,max}_{y\in\mathcal{N}_{\delta}(X^{\bullet})}f(y).$
Furthermore, we denote by $f(X^{\bullet})$ the common function value of $f$ on $X^{\bullet}$.
Figure 3: Illustration of a local maximizer set. The dotted unit circle is a local maximizer set of a real-analytic function and is not a singleton.
@@
Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be real-analytic. Assume $f$ has a global minimizer (not necessarily unique) and a bounded local maximizer set. Then, a hallucinated minimizer exists for some $\rho>0$.
The proof of Theorem 2.2, fully presented in Appendix A.4, is analogous to that of Theorem 2.1, except that the isolatedness assumption is replaced by the real-analyticity assumption. A key technical point is to rule out the possibility that critical points accumulate densely around $X^{\bullet}$. This is achieved using the Łojasiewicz inequality:
Lemma 2.3 (Łojasiewicz (1963)).
$|f(p)-f(x)|^{q}\leq C\|\nabla f(x)\|\quad\textit{ for all }\,x\in U.$
Most modern neural networks are real-analytic.
The real-analyticity assumption on $f$ is practical in the context of deep learning. In particular, consider the empirical loss function
$f(\theta)=\sum_{i=1}^{N}\ell(h_{\theta}(X_{i}),Y_{i}),$
where $\theta\in\mathbb{R}^{d}$ denotes the neural network parameters. If the dataset is finite ($N<\infty$), $\ell(\cdot,\cdot)$ is real-analytic in its first argument (as is the case for most commonly used losses, such as cross-entropy), and $h_{\theta}$ is a neural network built from real-analytic activation functions (e.g., tanh, ELU, GELU, SiLU, swish), then $f(\theta)$ is real-analytic. More concretely, $h_{\theta}$ may be a finite composition of linear layers, convolution, attention, residual connections, layer normalization, batch normalization, real-analytic activation functions, the softmax function, average pooling, and dropout. However, $h_{\theta}$ cannot incorporate ReLU, leaky ReLU, or max-pooling, since these operations are non-smooth and therefore non-analytic.
Further discussion and details are provided in Appendix A.3.
3 Geometric and dynamical properties of hallucinated minimizers
In this section, we establish a finer geometric property of hallucinated minimizers as well as a dynamical property of the SAM iterates. Recall that we use the notation $u(x)=\nabla f(x)/\|\nabla f(x)\|$.
3.1 The set of hallucinated minimizers can have manifold structures
The following theorem establishes that when the set of true minimizers has an $m$-dimensional manifold structure, the set of hallucinated minimizers inherits the same geometric structure. In particular, this result explains why, in Figure 1, the set of SAM minimizers appears as a curve.
@@ 3.1 The set of hallucinated minimizers can have manifold structures
Suppose $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfies the assumptions of Theorem 2.2. Assume $\mathcal{M}\subseteq\mathrm{argmin}f$, where $\mathcal{M}\subseteq\mathbb{R}^{d}$ is a nonempty smooth $m$-dimensional manifold. Let $x_{h}$ be a hallucinated minimizer with a corresponding $\rho>0$ as constructed in the proof of Theorem 2.2. If $I+\rho\nabla u(x_{h})\in\mathbb{R}^{d\times d}$ is nonsingular, then the set of hallucinated minimizers contains a smooth manifold of dimension $m$.
@@
3.2 Hallucinated minimizers are attractors
We have established the existence of hallucinated minimizers, but does SAM actually converge to them? Recall that the SAM iteration is given by
$x_{k+1}=x_{k}-\eta_{k}\nabla f\!\left(x_{k}+\rho\,u(x_{k})\right),\qquad k=0,1,2,\dots,$
where $\eta_{k}>0$ denotes the step size.
The answer is yes—hallucinated minimizers can indeed be attractors of the SAM dynamics. Thus, the concern about hallucinated minimizers in neural network training is not merely hypothetical. In Section 4, we provide empirical evidence that SAM can converge to hallucinated minimizers. In this subsection, we theoretically establish local convergence to hallucinated minimizers.
@@ 3.2 Hallucinated minimizers are attractors
Suppose $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is real-analytic, and let $H\subset\mathbb{R}^{d}$ be a bounded, connected set of hallucinated minimizers of $f$ for a fixed perturbation radius $\rho>0$. Assume there exists $\delta>0$ such that the $\delta$-neighborhood of $H$ contains no minimizers of $f^{\mathrm{SAM}}$ other than those already in $H$. Assume further that every $x_{h}\in H$ satisfies
$1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{h})))>0,\,\,\textup{ where }\,\mathrm{Sym}(\nabla u(x_{h}))=\frac{1}{2}(\nabla u(x_{h})+\nabla u(x_{h})^{\top}).$
If the initialization $x_{0}$ is chosen sufficiently close to $H$, then there exists a sufficiently small fixed step size $\eta_{k}=\eta>0$ such that the SAM iterates converge to $H$, in the sense that $d(x_{k},H)\to 0$.
We defer the proof to Appendix B.2.
4 Empirical analyses in deep learning
In this section, we empirically validate our theory by analyzing hallucinated minimizers in deep learning. We show that SAM trajectories can, in practice, converge to hallucinated minimizers. We further demonstrate that a simple switching strategy can effectively prevent this convergence, providing a practical safeguard for SAM against this failure mode.
@@ 4 Empirical analyses in deep learning
Figure 4: Visualizations of $f$ and $f^{\mathrm{SAM}}$ around the hallucinated minimizer $x_{h}$. Plots (a) and (b) are taken on a 2-dimensional plane defined by $x_{h}$ and two random directions. These show that $x_{h}$ is indeed a minimizer of $f^{\mathrm{SAM}}$ but not a stationary point of $f$. Plot (c) depicts $f^{\mathrm{SAM}}$ on the 2-dimensional plane containing $x_{h}$, $x_{0}$, and $x_{N}$, where $x_{0}$ is a small perturbation of $x_{h}$ and $x_{N}$ is obtained after $N=1000$ SAM steps from $x_{0}$. The pink horizontal line segment indicates the set of hallucinated minimizers, showing that the SAM trajectory converges back to this set.
4.1 SAM can converge to hallucinated minimizers
We first present an example in which SAM converges to a non-minimizer and confirm that the point is a hallucinated minimizer. We then examine how frequently such convergence occurs across diverse experimental settings.
Do we really converge to hallucinated minimizers?
We begin with a simple neural network setting and provide an example where SAM converges to a hallucinated minimizer. Specifically, we train a two-layer neural network with Tanh activations on MNIST (LeCun et al., 1998) using full-batch updates, yielding a smooth objective. The experimental setup is described in Appendix C. With a perturbation radius $\rho=1.8$, we train for over 20 million steps and observe that the trajectory converges to a single point $x_{h}$. At this point, the SAM gradient nearly vanishes ($\|\nabla f(x_{h}+\rho\,u(x_{h}))\|=4.8\times 10^{-9}$), while the original gradient remains relatively large ($\|\nabla f(x_{h})\|=0.0627$). Thus, $x_{h}$ is a stationary point of $f^{\mathrm{SAM}}$ but not of $f$, showing that SAM can converge to a hallucinated stationary point.
To verify that $x_{h}$ is indeed a hallucinated minimizer, we visualize the loss landscape around it, as shown in Figure 4. Following the method of Li et al. (2018), we define the plane
$x(\alpha,\beta)=x_{h}+\alpha u+\beta v,$
where $u,v\in\mathbb{R}^{d}$ are orthogonal vectors of equal norm. In Figure 4(a), $x_{h}$ is clearly not a minimizer of $f$, whereas in Figure 4(b), on the same plane, it appears as a minimizer of $f^{\mathrm{SAM}}$. This demonstrates that SAM can converge to a hallucinated minimizer in neural networks.
To investigate the geometry around $x_{h}$ in more detail, we add a small random perturbation to $x_{h}$, yielding a nearby point $x_{0}$ with $\|x_{0}-x_{h}\|=0.1$. Starting from $x_{0}$, we run $N=1000$ additional SAM steps to reach $x_{N}$. Figure 4(c) shows the SAM objective on the plane spanned by $x_{h}$, $x_{0}$, $x_{N}$. The SAM trajectory (blue) is projected onto this plane, while the SAM minimizers (pink) are computed within a tolerance of $10^{-9}$. We observe that hallucinated minimizers form a connected set, consistent with prior work on connected minimizers in neural networks (Garipov et al., 2018) and with Theorem 3.1, which guarantees that this structure is preserved.
@@ 4.1 SAM can converge to hallucinated minimizers
Figure 5: Training loss and test accuracy for SAM on MNIST with a 2-layer network using full-batch gradients. The top row shows the results from SAM-only training, while the bottom row shows the results from the switching strategy as described in Section 4.2. Final average test accuracies are 96.77%, 90.85%, 67.88%, and 13.36% for SAM-only training, compared to 97.77%, 97.73%, 97.69%, and 97.62% for the switching strategy. These results demonstrate that the switching strategy improves test accuracy and stabilizes outcomes across perturbation radii.
How common are hallucinated minimizers?
Next, we investigate convergence to hallucinated minimizers across multiple experiments, showing that this phenomenon can occur frequently in deep learning. Figure 5 presents SAM training outcomes in the same full-batch setting as before, with 80 distinct seeds and 100,000 iterations for perturbation radii $\rho=1.0,1.3,1.6,1.9$. The top row shows SAM-only results, while the bottom row corresponds to the switching strategy discussed later. In this subsection, however, we focus on the SAM-only setting.
The experiments reveal that convergence to hallucinated minimizers depends on the perturbation radius $\rho$. At $\rho=1.0$, training consistently converges to zero loss, indicating convergence to a true minimizer of the original loss. At $\rho=1.3$ and $1.6$, some trajectories stabilize at nonzero-loss points. At $\rho=1.9$, most trajectories converge to such nonzero-loss points, indicating that SAM predominantly reaches hallucinated minimizers.
These findings provide strong evidence that hallucinated minimizers are not rare anomalies but occur consistently in deep learning when the perturbation radius $\rho$ is large. This observation aligns with Theorem 2.1, whose proof requires the perturbation radius $\rho$ to exceed the distance between a minimizer and a maximizer. At the same time, even for the same $\rho$, trajectories may converge either to a hallucinated minimizer or to a true minimizer, consistent with Theorem 3.2, which depends on the initialization’s proximity to a hallucinated minimizer.
In the stochastic case, we train ResNet-18 (He et al., 2016) on CIFAR-100 (Krizhevsky, 2009) with mini-batch updates, following the FSAM implementation of Li et al. (2024). We observe a similar trend: larger perturbation radii lead to unstable training. Further details of the experimental settings and results are provided in Appendix C.

### v2

@@ 2.1 Simplified existence proof with isolated maximizers
We provide a brief sketch of the argument, with full details deferred to Appendix A.1. Figure 2 illustrates the key idea.
Let $\varepsilon>0$ and define $C_{\varepsilon}$ as the $(f(x^{\bullet})-\varepsilon)$-superlevel set restricted to a neighborhood of the isolated local maximizer $x^{\bullet}$. For sufficiently small $\varepsilon>0$, the set $C_{\varepsilon}$ is compact and satisfies: $(i)$ $f(x^{\bullet})-\varepsilon\leq f(x)\leq f(x^{\bullet})$ for all $x\in C_{\varepsilon}$; $(ii)$ $\nabla f$ does not vanish on $C_{\varepsilon}\setminus\{x^{\bullet}\}$; and $(iii)$ $f(x)=f(x^{\bullet})-\varepsilon$ for all $x\in\partial C_{\varepsilon}$.
Next, consider $g(x)=\|x-x_{\star}\|^{2}$ and choose
$x_{h}\in\operatorname*{arg\,max}_{x\in C_{\varepsilon}}g(x).$
Then $x_{h}\in\partial C_{\varepsilon}$. Moreover, since $\nabla f(x_{h})\neq 0$ and $\partial C_{\varepsilon}$ locally coincides with the level set $\{x:\ f(x)=f(x^{\bullet})-\varepsilon\}$ near $x_{h}$, there exists a neighborhood $V$ of $x_{h}$ such that
$\Sigma:=\{x\in V:\ f(x)=f(x^{\bullet})-\varepsilon\}$
is an embedded $C^{1}$ hypersurface and $\Sigma=V\cap\partial C_{\varepsilon}$.
Since $x_{h}$ maximizes $g$ over $C_{\varepsilon}$ and lies on $\partial C_{\varepsilon}$, it is also a local maximizer of $g$ over $\Sigma$. By the method of Lagrange multipliers, there exists $\lambda\in\mathbb{R}$ such that
$\nabla g(x_{h})+\lambda\nabla f(x_{h})=0,$
equivalently, $2(x_{\star}-x_{h})=\lambda\nabla f(x_{h})$. Furthermore, one can show that $\lambda>0$ (see Lemma A.4 in Appendix A.1).
Let $\rho:=\|x_{\star}-x_{h}\|$. Taking norms gives $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$, and hence
$x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{h}^{+}.$
Therefore, $\nabla f(x_{h}^{+})=\nabla f(x_{\star})=0$, while $\nabla f(x_{h})\neq 0$ by construction. Moreover, $f^{\mathrm{SAM}}(x_{h})=f(x_{h}^{+})=f(x_{\star})=\min f$. Thus, $x_{h}$ is a (global) minimizer of $f^{\mathrm{SAM}}$ and is therefore a hallucinated minimizer. Finally, since $x^{\bullet}\in C_{\varepsilon}$ and $x_{h}$ maximizes $\|x-x_{\star}\|$ over $C_{\varepsilon}$, we have $\rho=\|x_{\star}-x_{h}\|\geq\|x_{\star}-x^{\bullet}\|$.
As $\varepsilon$ varies while satisfying (i)(i)$(iii)$, the induced radius $\rho(\varepsilon):=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|$ depends continuously on $\varepsilon$ and is strictly increasing (Lemma A.5 in Appendix A.1). This yields an interval of radii $\rho$ for which hallucinated minimizers exist. ∎
@@
Importantly, the existence of a hallucinated minimizer also holds when $x_{\star}$ is a local minimizer, provided that $f$ has a locally Lipschitz gradient; see Appendix A.2 for details.
When hallucinated minimizers arise.
@@
The proof of Theorem 2.2 reveals conditions under which hallucinated minimizers are likely to arise. Figure 2 illustrates the core mechanism: the construction enforces $x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{h}^{+}$. This means that at $x_{h}$, the gradient points directly toward $x_{\star}$. Near a minimizer, gradients typically point away from the minimizer, so such an alignment is hard to obtain locally. The presence of a nearby maximizer provides a region where gradients can align in the required direction, which explains why nonconvexity is essential. The construction is also inherently a large-$\rho$ phenomenon: it yields radii satisfying $\rho=\|x_{\star}-x_{h}\|\geq\|x_{\star}-x^{\bullet}\|$, and the hallucinated minimizer $x_{h}$ lies near the maximizer region (inside $C_{\varepsilon}$).
2.2 Existence with local maximizer sets
Theorem 2.2 assumes that the local maximizer is isolated. In many modern objectives, however, symmetries and flat directions can produce non-isolated maximizers (e.g., a plateau or a manifold of maximizers). We therefore extend the existence result to the case of a local maximizer set, i.e., a connected set that maximizes $f$ within a neighborhood.
Definition 2.3 (Local maximizer set).
@@ 2.2 Existence with local maximizer sets
Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be continuous. A nonempty connected set $X\subseteq\mathbb{R}^{d}$ is a local maximizer set of $f$ if there exists $\delta>0$ such that
$X=\operatorname*{arg\,max}_{y\in\mathcal{N}_{\delta}(X)}f(y).$
In this case, $f$ is constant on $X$, and we denote this common value by $f(X)$.
The following theorem shows that hallucinated minimizers can still arise when the maximizers are non-isolated.
@@
Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be real-analytic. Assume $f$ has a global minimizer $x_{\star}$ (not necessarily unique) and a bounded local maximizer set $X$. Then, there exists a nontrivial interval of radii $\rho$ for which hallucinated minimizers exist, and this interval contains values with $\rho\geq d(x_{\star},X)$.
The proof of Theorem 2.4 is provided in Appendix A.3.
What changes relative to the isolated case.
When $X$ is not isolated, critical points may in principle accumulate near $X$, which could invalidate the “regular level set” argument needed at the boundary $\partial C_{\varepsilon}$. Real-analyticity rules out such pathological accumulation: via the Łojasiewicz inequality, we show that any critical point sufficiently close to $X$ with function value near $f(X)$ must lie in $X$ itself (Lemma A.9 in Appendix A.3). This ensures $\nabla f\neq 0$ on $\partial C_{\varepsilon}$, enabling the same Lagrange-multiplier construction as in Theorem 2.2. As in the isolated case, varying $\varepsilon$ yields a strictly increasing, continuous radius map $\rho(\varepsilon)=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|$, hence an interval of admissible radii.
On real-analyticity of $f$.
The real-analyticity assumption is used to invoke the Łojasiewicz inequality and exclude pathological accumulations of critical points near $X$. For finite-dataset objectives, this assumption holds whenever both the loss and the network mapping are built from real-analytic primitives (e.g., Tanh/sigmoid/softplus/GELU activations with standard smooth losses); see Appendix A.4 for a self-contained sufficient condition. In particular, it applies directly to our full-batch MNIST experiment with Tanh activations (Section 4.1). While this excludes nonsmooth architectures such as ReLU and max-pooling, we empirically observe hallucinated minimizers even in ReLU-based networks (Section 4).
3 Dynamical and geometric properties of hallucinated minimizers
In this section, we establish a dynamical property of the SAM iterates as well as a finer geometric property of hallucinated minimizers.
3.1 Hallucinated minimizers can be attractors
Section 2 established that hallucinated minimizers can exist for suitable radii. We now show that this phenomenon is not merely an algebraic possibility: an isolated connected component of hallucinated minimizers can be locally attracting for the (full-batch) SAM update.
@@ 3.1 Hallucinated minimizers can be attractors
Suppose $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is real-analytic, and let $H\subset\mathbb{R}^{d}$ be a bounded, connected set of hallucinated minimizers (for $f$ at radius $\rho$). Assume there exists $\delta>0$ such that the $\delta$-neighborhood of $H$ contains no minimizers of $f^{\mathrm{SAM}}$ other than those already in $H$. Assume further that every $x_{h}\in H$ satisfies
$1+\rho\,\lambda_{\min}\!\bigl(\mathrm{Sym}(\nabla u(x_{h}))\bigr)>0.$
If the initialization $x_{0}$ is chosen sufficiently close to $H$, then there exists a sufficiently small fixed step size $\eta_{k}=\eta>0$ such that the SAM iterates (1) satisfy $d(x_{k},H)\to 0$.
@@
Interpreting the stability condition.
Unlike gradient descent on $f^{\mathrm{SAM}}$, the update (1) does not a priori guarantee descent of the surrogate loss $f^{\mathrm{SAM}}$. The condition $1+\rho\,\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{h})))>0$ ensures a local acute-angle relation between the surrogate gradient $\nabla f^{\mathrm{SAM}}(x)$ and the SAM gradient $\nabla f(x^{+})$. Indeed, for any $x$ with $\nabla f(x)\neq 0$ for which $u$ is differentiable, we have $\nabla f^{\mathrm{SAM}}(x)=\bigl(I+\rho\,\nabla u(x)\bigr)^{\top}\nabla f(x^{+})$. Therefore, with $g:=\nabla f(x^{+})$,
$\langle\nabla f^{\mathrm{SAM}}(x),\,g\rangle=g^{\top}\!\bigl(I+\rho\,\mathrm{Sym}(\nabla u(x))\bigr)g.$
Hence, on a neighborhood where $I+\rho\,\mathrm{Sym}(\nabla u(x))\succeq\gamma I$ for some $\gamma>0$, we have $\langle\nabla f^{\mathrm{SAM}}(x),g\rangle\geq\gamma\|g\|^{2}$. Combining this with $L$-smoothness of $f^{\mathrm{SAM}}$ on that neighborhood yields the descent estimate
$f^{\mathrm{SAM}}(x-\eta g)\leq f^{\mathrm{SAM}}(x)-\eta\Bigl(\gamma-\tfrac{L\eta}{2}\Bigr)\|g\|^{2},$
so sufficiently small steps decrease $f^{\mathrm{SAM}}$ unless $g=0$. The proof of Theorem 3.1 formalizes this argument near $H$ and uses the assumption that $H$ is an isolated minimizer component of $f^{\mathrm{SAM}}$ to conclude $d(x_{k},H)\to 0$.
A simple sufficient condition.
Directly verifying the eigenvalue condition in Theorem 3.1 can be difficult in high dimensions, so we record a convenient sufficient condition. At any $x$ with $\nabla f(x)\neq 0$ and $f\in C^{2}$ in a neighborhood of $x$, we have
$\nabla u(x)=\frac{1}{\|\nabla f(x)\|}\bigl(I-u(x)u(x)^{\top}\bigr)\nabla^{2}f(x),$
and thus $\|\nabla u(x)\|\leq\|\nabla^{2}f(x)\|/\|\nabla f(x)\|$. Using $\lambda_{\min}(\mathrm{Sym}(A))\geq-\|A\|$, the stability condition $1+\rho\,\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{h})))>0$ is implied by
$\rho\,\|\nabla^{2}f(x_{h})\|<\|\nabla f(x_{h})\|.$
This ratio-type condition is meaningful at hallucinated minimizers since $\|\nabla f(x_{h})\|>0$ by definition.
3.2 Hallucinated minimizers can form manifolds
We next refine the geometry of hallucinated minimizers. In the toy example of Figure 1, they are not isolated and instead form a curve-like set. The following result shows that this is not accidental: whenever the original objective has a manifold of global minimizers, the set of hallucinated minimizers can inherit the same manifold dimension.
@@ 3.2 Hallucinated minimizers can form manifolds
Assume $f:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{2}$. Let $\mathcal{M}\subseteq\operatorname*{arg\,min}f$ be a nonempty embedded $m$-dimensional $C^{1}$ manifold, and fix $x_{\star}\in\mathcal{M}$. Suppose there exists a point $x_{h}\in\mathbb{R}^{d}$ such that
$x_{h}^{+}=x_{\star}\quad\text{and}\quad\nabla f(x_{h})\neq 0.$
If in addition $I+\rho\nabla u(x_{h})$ is nonsingular, then there exist neighborhoods $V$ of $x_{h}$ and $U$ of $x_{\star}$ such that the set
$\mathcal{H}:=\{x\in V:\ x^{+}\in\mathcal{M}\cap U\}$
is an embedded $m$-dimensional $C^{1}$ manifold. Moreover, every $x\in\mathcal{H}$ is a hallucinated minimizer at radius $\rho$.
Its proof is provided in Appendix B.2.
Interpretation.
The nonsingularity of $I+\rho\nabla u(x_{h})$ exactly means that the perturbation map $\Phi_{\rho}(x):=x+\rho u(x)$ is locally invertible at $x_{h}$ (inverse function theorem). Thus, near $x_{h}$, the preimage $\Phi_{\rho}^{-1}(\mathcal{M})$ is a smooth manifold of the same dimension as $\mathcal{M}$.
Connection to our existence results.
Under the assumptions of Theorem 2.4, we may choose $x_{\star}$ to lie on any manifold component $\mathcal{M}\subseteq\operatorname*{arg\,min}f$ (when such a component exists). The construction in Section 2 then produces an $x_{h}$ with $x_{h}^{+}=x_{\star}$, and Theorem 3.2 explains why the resulting hallucinated minimizers can appear as curve-/surface-like sets in practice (e.g., Figure 1 and Figure 3).
4 Experiments
@@
Figure 3: Visualizations of $f$ and $f^{\mathrm{SAM}}$ around the hallucinated minimizer $x_{h}$. Plots (a) and (b) are taken on a two-dimensional plane defined by $x_{h}$ and two random directions. These show that $x_{h}$ is not stationary for $f$, while it appears as a minimizer of $f^{\mathrm{SAM}}$ on the same plane. Plot (c) depicts $f^{\mathrm{SAM}}$ on the two-dimensional plane containing $x_{h}$, $x_{0}$, and $x_{N}$, where $x_{0}$ is a small perturbation of $x_{h}$ and $x_{N}$ is obtained after $N=1000$ SAM steps from $x_{0}$. The pink horizontal line segment indicates the set of hallucinated minimizers, showing that the SAM trajectory converges back to this set.
We empirically validate the mechanisms identified in Sections 23 on neural network objectives. The implementation of our experiments is available through a repository at https://anonymous.4open.science/r/SAM-can-hallucinate-minimizers-4B82/.
Diagnostics.
To diagnose convergence to hallucinated minimizers along a trajectory $\{x_{k}\}$, we monitor two pointwise mismatch quantities $\texttt{GradRatio}(x):=\frac{\|\nabla f(x^{+})\|}{\|\nabla f(x)\|+\epsilon}$ and $\texttt{LossGap}(x):=f(x^{+})-f(x)$, with $\epsilon>0$ a small numerical stabilizer. At a hallucinated minimizer $x_{h}$, we expect $\|\nabla f(x_{h}^{+})\|\approx 0$ while $\|\nabla f(x_{h})\|>0$, hence $\texttt{GradRatio}(x_{h})\ll 1$. Moreover, since $x_{h}^{+}$ lies near a (local) minimizer of $f$, the “ascent” perturbation can paradoxically decrease the loss, yielding $\texttt{LossGap}(x_{h})\ll 0$. We emphasize that in neural networks these diagnostics are evidence consistent with hallucinated minimizers, rather than a formal certification.
@@
[…]
4.1 SAM can converge to hallucinated minimizers
We first provide a direct verification in a smooth, full-batch setting, and then consider larger networks and stochastic settings.
Direct verification in a smooth full-batch setting.
We train a two-layer network with Tanh activations on MNIST using full-batch SAM. This setting yields a smooth (indeed, real-analytic) objective, aligning with the assumptions underlying our existence and attractor results. Implementation details are in Appendix C.1. With perturbation radius $\rho=1.8$, we run training for a long horizon until the trajectory numerically stabilizes near a point $x_{h}$. At this point, the shifted gradient is nearly zero while the original gradient is not: $\|\nabla f(x_{h}^{+})\|=4.8\times 10^{-9}$, and $\|\nabla f(x_{h})\|=6.27\times 10^{-2}$. Thus, the practical SAM update becomes (nearly) stationary at $x_{h}$ even though $x_{h}$ is not a critical point of $f$.
To further probe the local geometry, we visualize two-dimensional loss landscapes around $x_{h}$ following Li et al. (2018). Specifically, we consider planes of the form $x(\alpha,\beta)=x_{h}+\alpha d_{1}+\beta d_{2}$ where $d_{1},d_{2}$ are random orthogonal directions of matched norm. Figure 3 shows that $x_{h}$ is not a minimizer of $f$ on this plane, while it appears as a local minimizer of the SAM loss $f^{\mathrm{SAM}}(x)=f(x^{+})$ on the same slice. This is consistent with $x_{h}$ behaving as a hallucinated minimizer.
Local attraction in practice.
To test whether the observed point is locally attracting (rather than an artifact of a single run), we perturb $x_{h}$ by a random displacement of size $0.1$ to form $x_{0}$, run $N=1000$ additional SAM steps, and examine the resulting point $x_{N}$. Figure 3(c) shows that the perturbed initialization returns to the same near-minimizer region of $f^{\mathrm{SAM}}$ around $x_{h}$ within this local slice, suggesting robust local attraction, consistent with Theorem 3.1. Moreover, the near-minimizer region is visibly non-isolated on the slice, qualitatively aligning with the manifold mechanism in Theorem 3.2 (see also Appendix C.1).
How common is the failure mode? (full-batch MNIST)
We next sweep $\rho\in\{1.0,1.3,1.6,1.9\}$ and run 80 independent seeds for 100,000 iterations. Figure 4 reports training loss and test accuracy. For $\rho=1.0$, trajectories consistently reach near-zero loss. As $\rho$ increases, a substantial fraction of runs plateau at positive training loss and exhibit severe accuracy degradation, matching the trajectory-level signature of convergence to a spurious stationary region for the practical SAM update. Notably, at the same $\rho$, different seeds can yield either near-zero loss or a plateau, consistent with the basin-of-attraction viewpoint in Theorem 3.1.
Table 1: SAM with full-batch gradients on CIFAR-10/100 (ResNet-18): metrics at the final iterate $x_{T}$ across perturbation radii $\rho$. Here, $x_{T}^{+}=x_{T}+\rho\,u(x_{T})$; $f(x_{T}^{+})$ is the SAM loss; and $\nabla f(x_{T}^{+})$ is the SAM gradient.
CIFAR-10 CIFAR-100
$\rho$ $f(x_{T})$ $f(x_{T}^{+})$ $\lVert\nabla f(x_{T})\rVert$ $\lVert\nabla f(x_{T}^{+})\rVert$ $f(x_{T})$ $f(x_{T}^{+})$ $\lVert\nabla f(x_{T})\rVert$ $\lVert\nabla f(x_{T}^{+})\rVert$
0.0 0.001 0.001 0.006 0.006 0.009 0.009 0.058 0.058
0.5 0.081 0.001 3.190 0.015 0.050 0.009 0.944 0.100
1.0 0.120 0.001 3.131 0.015 0.132 0.011 4.034 0.134
1.5 0.773 0.001 11.932 0.011 0.419 0.012 8.119 0.136
2.0 0.219 0.002 4.033 0.021 0.828 0.014 13.722 0.123